# Agent-as-Judge for Factual Summarization of Long Narratives

### v1

@@ 5.2 Correlation with Human Factuality Scores
     Agent Refinement, 1st round (Ours) 33.00 9.70 12.84 60.22 60.11 60.16 85.94 30.36 8.74 18.55 59.26 59.30 59.27 86.92
     Agent Refinement, 2nd round (Ours) 33.75 9.72 13.07 60.17 60.10 60.12 88.94 30.98 8.75 18.61 59.33 59.30 59.30 92.04
     Agent Refinement, 3rd round (Ours) 34.26 9.74 13.46 60.24 60.21 60.22 93.31 31.31 8.81 18.62 59.36 59.31 59.32 94.59
@@
Dataset
To evaluate whether the NarrativeFactScore we proposed correlates effectively with human factuality, we conducted a series of experiments. For this purpose, we used STORYSUMM  Subbiah et al. (2024) and FABLES Kim et al. (2024), which include multiple summaries generated by LLM for each narrative.
Correlation with Human Annotations
These summaries were then evaluated by human annotators based on their factual accuracy. We computed the Spearman Spearman (1961) correlations and Kendall’s tau KENDALL (1938) correlations for each metric in relation to the human factuality scores, as shown in Table 1.
@@ Results
As observed in the results, NarrativeFactScore shows the highest correlation with human annotations among all metrics. Specifically, NarrativeFactScore achieves a Kendall’s tau correlation coefficient exceeding 0.3444When interpreting Kendall’s $\tau$, $|\tau|\in[0,0.1)$ is considered as very weak correlation, $|\tau|\in[0.1,0.2)$ is considered as weak correlation, $|\tau|\in[0.2,0.3)$ is considered as moderate correlation, and $|\tau|\in[0.3,1.0]$ is considered as strong correlation Chiang and Lee (2023)., which indicates a strong correlation with human annotations. The statistical significance of this correlation is calculated with a p-value of 0.00003, indicating that the relationship between NarrativeFactScore and human annotations is strong and statistically significant.
5.3 Summarization Performance Evaluation
Datasets
We evaluated our framework on the MENSA Saxena and Keller (2024b) and MovieSum Saxena and Keller (2024a) datasets. MENSA aligns movie scenes with Wikipedia summaries, and MovieSum pairs screenplays with summaries. We use the full test sets: 50 samples from MENSA and 200 from MovieSum.
Baselines
@@ 5.3 Summarization Performance Evaluation
We evaluated summarization performance using two baseline types. The first type includes methods without merging that summarize all input in a single step, such as TextRank Mihalcea and Tarau (2004), Longformer Encoder-Decoder (LED) Beltagy et al. (2020), and LongT5 Guo et al. (2022). The second type involves hierarchical merging Chang et al. (2024), with which we performed experiments using GPT-4o-mini OpenAI (2023). Additionally, we evaluated the summaries generated by GPT-4o-mini after iterative improvements (1st to 3rd refinements) through agent-based refinement.
Results
As shown in Table 2, agent-based refinement improves not only factuality but also the evaluation metrics across the board, improving the overall quality of the summaries. The results from the non-merging baselines indicate that summarizing long narratives is challenging both in terms of factuality and other metrics. However, agent-based refinement demonstrates the ability to improve performance, although repeated refinements may lead to saturation, where further improvements are minimal. Nevertheless, since our refinement primarily targets factuality, improvements in factuality remain consistent even after multiple refinements.
6 Conclusion
This demonstration shows how the agent-as-judge contributes to overcoming the limitations of existing evaluation metrics, such as overreliance on lexical similarity or factual inconsistencies. Specifically, we propose consistent CKG extraction, and new factual evaluation metric based on CKG, and a agent that evaluates and guides the summary and refinement. Through our implementation, we demonstrated both the process and superior performance over state-of-the-art methods on real-life industry datasets and scenarios.
Appendices
Appendix A Analysis
A.1 Analysis for Baseline Metrics
@@
We investigate the interpretation of the results compared to other metrics in Table 1. Metrics based on lexical overlap, such as ROUGE, show stronger correlations with human factuality assessments compared to semantic similarity metrics such as BERTScore, as they better capture repeated entities and locations in narratives. In contrast, metrics such as BARTScore and LongDocFACTScore Bishop et al. (2024), which rely on log-likelihood and entailment, have lower correlations due to their limited ability to account for broader context and character relationships. FActScore Min et al. (2023), reproduced in our study, incorporates character relationship retrieval to improve factuality assessments. Building on this, NarrativeFactScore further enhances performance by addressing common errors caused by misinterpreted character relationships, leading to more accurate evaluations.
A.2 How Consistently Does Ours Capture Character Relationships?
@@
To effectively evaluate factuality and improve summary, it is necessary to generate an accurate and consistent CKG. According to Kim et al. (2024); Zhao et al. (2024), the “naive extract” approach, where an LLM extracts character relationships from a story in one step, often fails to consistently capture some relationships. Thus, our objective is to verify whether our approach can generate a consistent CKG. Conan Zhao et al. (2024) provides ground truth annotation of character relationships within narratives. To evaluate whether the generated relation is semantically similar to this ground truth, we measure the BERTScore Devlin et al. (2019).
@@ A.3 Challenging Set
     Agent Refinement, 1st round (Ours) 29.02 8.09 14.08 58.49 59.30 58.86 84.59
     Agent Refinement, 2nd round (Ours) 29.98 8.19 14.24 58.61 59.32 58.94 90.47
     Agent Refinement, 3rd round (Ours) 30.22 8.20 14.44 58.75 59.39 59.04 93.22
@@
Our metric demonstrates the capability to provide feedback for improving factuality even in recent works. For this experiment, we curated a challenging set of 18 movies from MovieSum Saxena and Keller (2024a) released after our LLM knowledge cutoff date  [^5] We used GPT-4o-mini with an October 2023 knowledge cutoff date.. We conducted refinement experiments identical to Table 2 to improve the factual inaccuracies in this challenging set. As shown in Table 4, three rounds of refinement improved NarrativeFactScore by 11.92, comparable to the improvements in Table 2. These results confirm our metric provides effective factuality-improving feedback for recent works independent of LLM parametric knowledge.
Appendix B Implementation Details
Uniform Language Model Usage
To ensure that performance improvements are due to our framework and not the underlying language model, we use only gpt-4o-mini-2024-07-18 in our experiments. This model is applied across all components, including CKG extraction, summarization, fact decomposition, fact check, and agent-based fact refinement. This approach prevents superior LLMs from influencing the results, allowing us to rigorously evaluate the effectiveness of our framework.
Generating Initial Summary
To generate the initial summary $Z$, we adopt hierarchical merging Chang et al. (2024) that ensures the logical structure of the narrative is preserved. The narrative is first divided into chunks $C_{i}$ where each chunk is formed incrementally by adding scenes until a predefined context size666we set predefined context size of a chunk to 1024 is reached. Once this limit is exceeded, a new chunk begins, resulting in a sequence of chunks $\mathcal{C}=\{C_{1},C_{2},...,C_{n}\}$. Each chunk $C_{i}$ is then independently summarized using the prompt specified in Section D.2, and the resulting chunk summaries are sequentially merged to produce the initial summary $Z$.
Retrieving Relevant Scene and Subgraph
Using the BGE-M3 embedding model Chen et al. (2024), we retrieve information relevant to each atomic fact $a_{k}$. Specifically, we identify the most similar scene $\mathcal{S}_{i}$ from the narrative $\mathcal{N}$ and a subgraph containing the three most relevant triples in the linearized CKG $G$.
Appendix C Demonstration Example
Figure 4: Demonstration of summarizing narrative.

### v2

@@ 4.4 Summarization Performance Evaluation
     Ours: 1st iteration 33.00 9.70 12.84 60.22 60.11 60.16 85.94 30.36 8.74 18.55 59.26 59.30 59.27 86.92
     Ours: 2nd iteration 33.75 9.72 13.07 60.17 60.10 60.12 88.94 30.98 8.75 18.61 59.33 59.30 59.30 92.04
     Ours: 3rd iteration 34.26 9.74 13.46 60.24 60.21 60.22 93.31 31.31 8.81 18.62 59.36 59.31 59.32 94.59
@@
Datasets
We evaluated our framework on the MENSA Saxena and Keller (2024b) and MovieSum Saxena and Keller (2024a) datasets. MENSA aligns movie scenes with Wikipedia summaries and MovieSum pairs screenplays with summaries. We use the full test sets: 50 samples from MENSA and 200 from MovieSum.
@@ Results
We evaluated summarization performance using two baseline types. The first type includes methods without merging that summarize all input in a single step, such as TextRank Mihalcea and Tarau (2004), Longformer Encoder-Decoder (LED) Beltagy et al. (2020), and LongT5 Guo et al. (2022). The second type involves hierarchical merging Chang et al. (2024), with which we performed experiments using GPT-4o-mini OpenAI (2023). Additionally, we evaluated the summaries generated by GPT-4o-mini after agent-based iterative refinements (1st to 3rd).
As shown in Table 3, agent-based refinement improves not only factuality but also other metrics, improving the overall quality of the summaries. This refinement improves performance consistently, yielding improvements of +14.03 in factuality, +2.05 in ROUGE, and +0.13 in BERTScore on MovieSum, and +12.26, +2.47, and +0.21 respectively on MENSA.
5 Analysis
5.1 Analysis for Baseline Metrics
Strength Spearman ($\rho$) KENDALL ($\tau$)
Very weak correlation 0.00~0.15 0.00~0.10
Weak correlation 0.15~0.30 0.10~0.20
Moderate correlation 0.30~0.43 0.20~0.30
Strong correlation 0.43~1.00 0.30~1.00
Table 4: Correlation strength based on coefficients.
[…]
@@ 5.1 Analysis for Baseline Metrics
We analyze the results compared to other metrics in Table 6. Metrics based on lexical overlap, such as ROUGE, show stronger correlations with human factuality assessments compared to semantic similarity metrics such as BERTScore, as they better capture repeated entities and locations in narratives. In contrast, metrics such as BARTScore and LongDocFACTScore Bishop et al. (2024), which rely on log-likelihood and entailment, have lower correlations due to their limited ability to account for broader context and character relationships. FActScore Min et al. (2023), reproduced in our study, incorporates character relationship retrieval to improve factuality assessments. Building on this, NarrativeFactScore further enhances performance by addressing common errors caused by misinterpreted character relationships, leading to more accurate evaluations.
5.2 How Consistently Does Ours Capture Character Relationships?
@@
To effectively evaluate factuality and improve summary, it is necessary to generate an accurate and consistent CKG. According to Kim et al. (2024); Zhao et al. (2024), the “naive extract” approach, where an LLM extracts character relationships in one step, often fails to consistently capture some relationships. Thus, our objective is to verify whether our approach can generate a consistent CKG. Conan Zhao et al. (2024) provides ground truth annotation of character relationships within narratives. To evaluate whether the generated relation is semantically similar to this ground truth, we measure the BERTScore Devlin et al. (2019).
@@ 5.3 Challenging Set
     Ours: 1st iteration 29.02 8.09 14.08 58.49 59.30 58.86 84.59
     Ours: 2nd iteration 29.98 8.19 14.24 58.61 59.32 58.94 90.47
     Ours: 3rd iteration 30.22 8.20 14.44 58.75 59.39 59.04 93.22
@@
Our metric demonstrates the capability to provide feedback for improving factuality even in recent works. For this experiment, we curated a challenging set of 18 movies from MovieSum Saxena and Keller (2024a) released after our LLM knowledge cutoff. [^8] We used GPT-4o-mini with an October 2023 knowledge cutoff date. We conducted refinement experiments identical to Table 3 to correct factual errors in this challenging set. As shown in Table 6, three rounds of refinement improved NarrativeFactScore by 11.92, comparable to the improvements in Table 3. These results confirm our metric provides effective feedback for recent stories independent of LLM parametric knowledge.
5.4 Latency
Metric
CKG
extraction
time
Factuality
calculation
time
KENDALL
Human - 132.00 min -
LongDocFACTScore - 3.81 min 0.16
FActScore - 4.60 min 0.09
NarrativeFactScore 1.17 min 4.81 min 0.33
Table 7: Average latency (in minutes) and KENDALL’s tau correlation for evaluating factuality across different metrics on the FABLES Kim et al. (2024) dataset.
We compared the latency of LLM-as-a-Judge metrics, such as LongDocFACTScore and FActScore, with our Agent-as-a-Judge metric. Table 7 shows the time required to evaluate the factuality of each summary in the FABLES Kim et al. (2024). Since long narratives like those in FABLES exceed 100K tokens, human evaluation by verifying details is time-consuming. [^9] According to Kim et al. (2024), annotators spent over 11 hours evaluating five summaries. In contrast, LLM-based metrics, including ours, assess factuality within a few minutes. However, LLM-as-a-Judge metrics struggle to assess factuality while understanding character relationships, leading to discrepancies with human evaluations, as shown in Table 6. In contrast, NarrativeFactScore devotes additional time to reasoning about character relationships before assessing factuality, resulting in more accurate evaluations despite slightly longer times.
5.5 Sensitivity of Metrics on Factual Perturbation
Summary ROUGE-L BERTScoref1 NFS
Reference Summary 100.00 100.00 95.42
Perturbed Summary 81.61 92.15 40.81
Table 8: Change in metric scores after factual perturbation of the reference summary on the MENSA.
We evaluated metric sensitivity to factual perturbations using GPT-4o with the prompt shown in Figure 10 on the MENSA. Specifically, we perturbed the reference summaries from MENSA by introducing factual inaccuracies in each sentence. Table 8 shows that ROUGE-L and BERTScoref1 decreased minimally despite factual perturbations, while NarrativeFactScore significantly dropped. These results demonstrate that our metric is highly sensitive to factual discrepancies, making it a suitable metric for assessing factuality.
6 Conclusion
This work shows how the agent-as-judge is deployed to overcome the limitations of existing evaluation metrics, such as overreliance on lexical similarity or factual inconsistencies. Specifically, we propose consistent CKG extraction, and new factual evaluation metric based on CKG, and an agent that evaluates and guides the summary and refinement. Through our implementation, we demonstrated both the process and superior performance over state-of-the-art methods on real-life industry datasets and scenarios.
7 Limitation
We acknowledge two limitations of our work. First, our framework may occasionally retrieve subgraphs unrelated to the atomic fact being evaluated, though this did not impact factuality judgments in our experiments and outperformed the no-retrieval baseline. Nonetheless, further enhancing subgraph retrieval precision remains a promising direction for future work.
Second, our framework has been tested exclusively in the narrative domain. Although effective, its generalizability to other domains remains unverified. However, its potential for applications requiring deep character understanding—such as news summarization, biographical writing, and historical analysis—suggests promising directions for future exploration.
Appendices
Appendix A Qualitative Example
Initial
summary
Saruman reveals to Gandalf that Sauron is regaining strength and is gathering an army to attack Middle-earth, using a Palantir to see his plans. Gandalf warns Saruman about the dangers of the Palantir and expresses disbelief at Saruman’s willingness to ally with Sauron, leading to a fierce battle between the two. In the subsequent scene, Frodo and Sam are enjoying a peaceful moment in the Shire when they are interrupted by Merry and Pippin, who crash into them after stealing vegetables from Farmer Maggot’s field, prompting a humorous chase as they flee from the angry farmer and his dogs. The outcome sees Gandalf and Saruman in conflict, while Frodo and his friends are caught up in a lighthearted escapade.
Atomic
facts
[1] Saruman reveals to Gandalf that Sauron is regaining strength.
[2] Sauron is gathering an army to attack Middle-earth.
[3] Sauron uses a Palantir to see his plans.
[4] Gandalf warns Saruman about the dangers of the Palantir.
[5] Gandalf expresses disbelief at Saruman’s willingness to ally with Sauron.
[6] A fierce battle occurs between Gandalf and Saruman.
[7] Frodo and Sam enjoy a peaceful moment in the Shire.
[8] Merry and Pippin crash into Frodo and Sam.
[9] Merry and Pippin steal vegetables from Farmer Maggot’s field.
[10] A humorous chase ensues as they flee from Farmer Maggot and his dogs.
[11] Gandalf and Saruman are in conflict.
[12] Frodo and his friends are caught up in a lighthearted escapade.
Factuality
calculation
[1] True
[2] True
[3] False, The statement is false because Sauron does not use a Palantir to see his plans; rather, it is Saruman who uses the Palantir to gain knowledge about Sauron’s actions and intentions.
(evidence scene: #39, evidence subgraph: ‘Saruman-own-Palantir’, …)
[4] True
[5] True
[6] True
[7] False, The statement “Frodo and Sam enjoy a peaceful moment in the Shire” is false. The scene depicts Frodo and Sam being interrupted by Merry and Pippin, leading to a messy situation as they are chased by Farmer Maggot and his dogs after stealing vegetables from his field.
(evidence scene: #40, evidence subgraph: ‘Frodo-friend-Sam’, …)
[8] True
[9] True
[10] True
[11] True
[12] True
Refined
summary
Saruman reveals to Gandalf that Sauron is regaining strength and is gathering an army to attack Middle-earth. Saruman has used a Palantir to gain insight into Sauron’s plans. Gandalf warns Saruman about the dangers of the Palantir and expresses disbelief at Saruman’s willingness to ally with Sauron, leading to a fierce battle between the two. In the subsequent scene, Frodo and Sam are caught in a messy situation in the Shire when Merry and Pippin crash into them after stealing vegetables from Farmer Maggot’s field, prompting a humorous chase as they flee from the angry farmer and his dogs. The outcome sees Gandalf and Saruman in conflict, while Frodo and his friends are caught up in a lighthearted escapade.
Table 9: Qualitative example illustrating NarrativeFactScore. Red text in the initial summary and atomic facts indicate factually incorrect statements based on scene evidence, while blue text in the refined summary indicate corrections made through agent-based refinement based on feedback.
In this section, we illustrate how our approach evaluates factuality, provides actionable feedback, and refines summaries through a qualitative example. Table 9 shows the evaluation and refinement process for a summary of The Lord of the Rings generated by GPT-4o-mini OpenAI (2023). The CKG has already been constructed using the method described in Section 3.1.
To evaluate factuality, we decompose the initial summary into atomic facts. The summary contains two factually incorrect statements highlighted in red, which are also presented in the atomic facts. First, according to the original script, Saruman uses a Palantir to observe Sauron; however, due to difficulty in understanding character relationships, the incorrect summary stating Sauron observes Saruman was generated and recorded as atomic fact [3]. Second, while the original script depicts Frodo and Sam in a messy situation, chased in the Shire by Merry and Pippin, the summary incorrectly describes it as peaceful, recorded as atomic fact [7].
In our framework, we retrieve relevant scene and subgraph for each atomic fact to evaluate factuality. Consequently, only [3] and [7] were identified as false. For these facts, the framework generates not only the factuality but also actionable feedback explaining why they are false and how to correct them. For [3], based on retrieved scene and relationship that Saruman owns the Palantir, our framework determines that the statement is false and generates feedback suggesting that Saruman uses the Palantir to gain knowledge about Sauron’s actions and intentions. Similarly, for [7], based on the scene showing Frodo and Sam in a messy situation with Merry and Pippin in the Shire, our framework determines the statement is false and provides proper feedback. Using this detailed feedback, our framework generates refined summary by correcting only the erroneous parts of the initial summary.
Appendix B System Deployment
Figure 4: Deployment overview of NarrativeFactScore.
This section describes our system deployment, which is necessary for the media industry [^10] Company name anonymized for blind reviewing. The media industry broadly refers to the sector that creates, distributes, and analyzes various forms of narratives, such as movies, television shows, books, video games, and other media that tell stories. This includes businesses involved in producing, editing, and consuming these forms of content, focusing on storytelling in both traditional and digital media. where companies make investment decisions on narratives (> 100K tokens) such as dramas or movies. Since reading all narratives is challenging, media companies utilize summaries of each narrative to determine its production feasibility. However, summaries generated by humans or LLMs frequently contain factual inconsistencies, which hinder accurate investment decisions. Therefore, our proposed system is deployed to evaluate the factuality of summary for long narratives and improve its factuality.
Figure 4 shows screenshots of the system111111bit.ly/4iz8pXc, aligned with the three phases of our framework in Figure 2. Using the example Black Panther, users can view the original narrative after selecting a dataset, data type, and name. Clicking “Generate Knowledge Graph” generates and visualizes the CKG (Section 3.1). The “Generate Initial Summary” and “Calculate Factuality Score” buttons create an initial summary and evaluate its factuality using the CKG (Section 3.2). Finally, “Refine Summary” improves the summary based on feedback, enhancing factuality (Section 3.3).
Appendix C Usage of AI Assistants
We utilized ChatGPT to improve the clarity and grammatical accuracy of my writing. It provided suggestions for rephrasing sentences and correcting grammatical errors to make the text flow more naturally.
@@ D.5 Agent-based Refinement Prompt for LLM
D.6 Factual Perturbation Prompt
Narrative Summarization Prompt This sentence serves as a summary of a script. Rewrite this one-sentence summary by minimally replacing a few words in the original sentence to render it factually inaccurate, while keeping the original sentence structure intact. Original sentence: {original_sentence} Rewritten sentence:
Figure 10: Prompt used to generate factual perturbations.